OpenStudy (joclyn16):

The surface area of a cylinder is 80,603.8 square millimeters. The radius is 85. What is the height?

OpenStudy (danjs):

a cylinder is composed of 2 circles and the rectangle side wall the side wall dimensions are , height by circumference of circle

OpenStudy (danjs):

\[A = \pi*r^2 + \pi*r^2 + 2\pi*r*h\] there it is, two same circles, and the side, circumference by h

OpenStudy (danjs):

solve the height, A=80603.8 and r=85

OpenStudy (joclyn16):

So divide 80603.8/85

OpenStudy (joclyn16):

Help

OpenStudy (joclyn16):

Now

OpenStudy (joclyn16):

I just need the answer

OpenStudy (mathstudent55):

Copy @DanJS 's formula. Replace r with the radius. Replace A with the area. The only unknown left is h. Solve for h.

OpenStudy (joclyn16):

I don't get it

OpenStudy (mathstudent55):

Do you know what "copy" means, and do you know what "replace" means?

OpenStudy (joclyn16):

Yes

OpenStudy (mathstudent55):

here is the formula: \(A = \pi*r^2 + \pi*r^2 + 2\pi*r*h\) Write this formula, but write 85 instead of r, and write 80,603.8 instead of A.

OpenStudy (joclyn16):

But I don't know what the height, h, is

OpenStudy (mathstudent55):

Now you should have an equation that has only one unknown, h.

OpenStudy (mathstudent55):

Can you write the equation with those two replacements. Then we will solve it for h together, and you'll find the height.

OpenStudy (joclyn16):

Ok

OpenStudy (mathstudent55):

Also, you can use 3.14 for pi, so you get this: \(80,603.8 = 3.14 \times (85)^2 + 3.14 \times (85)^2 + 2 \times 3.14 \times 85 \times h\)

OpenStudy (mathstudent55):

Now we have an equation. The equation has only one variable, h. We can solve for h.

OpenStudy (joclyn16):

Ok

OpenStudy (joclyn16):

45,906.8

OpenStudy (mathstudent55):

Let's start working on the numbers on the right side. \(80,603.8 = 3.14 \times \color{red}{(85)^2} + 3.14 \times \color{red}{(85)^2} + 2 \times 3.14 \times 85 \times h\) What is \(85^2\) ?

OpenStudy (joclyn16):

7,225

OpenStudy (mathstudent55):

Good. \(80,603.8 = \color{red}{3.14 \times 7225} + \color{red}{3.14 \times 7225} + 2 \times 3.14 \times 85 \times h\) What is 3.14 * 7225 ?

OpenStudy (joclyn16):

22,686.5

OpenStudy (mathstudent55):

Good. \(80,603.8 = \color{red}{26,686.5 + 22,686.5} + 2 \times 3.14 \times 85 \times h\) Can you add the numbers in red?

OpenStudy (joclyn16):

49,373.3

OpenStudy (mathstudent55):

Sorry, I had a typo above. I meant 22,686.5 + 22,686.5 The sum is 45,373 \(80,603.8 = \color{red}{45,373} + 2 \times 3.14 \times 85 \times h\) Now we subtract 45,373 from, both sides. What is 80,603.3 - 45,373 ?

OpenStudy (joclyn16):

35,230.3

OpenStudy (mathstudent55):

You did it correctly. Sorry, I had another typo. It was 80,603.8 - 45,373 = 35,230.8 Now we have: \(35,230.8 = \color{red}{2 \times 3.14 \times 85} \times h\) Now multiply together the numbers in red.

OpenStudy (joclyn16):

533.8

OpenStudy (mathstudent55):

Good. \(35,230.8 = 533.8 \times h\) Now we need to divide both sides by 533.8. What is 35,230.8 / 533.8 ?

OpenStudy (joclyn16):

66

OpenStudy (mathstudent55):

Good. We now have 66 = h The height is 66 mm.

OpenStudy (joclyn16):

Thanks

OpenStudy (mathstudent55):

You're welcome.